Ten grams of Uranium will decay to 2.5 g in 496,000 years. What is its half-life?

Answer by lwsshak3(11628)

Ten grams of Uranium will decay to 2.5 g in 496,000 years. What is its half-life?
**
Formula for radioactive decay: A/Ao=2^-t/h, A=radioactive material present at time t, Ao=amt present initially, h=material half life.
2.5/10=2^-4.96*10^5/h
.25=2^-4.96*10^5/h
take log of both sides
log.25=-4.96*10^5/h)log2
h(log.25)=-(4.96*10^5)log2
h=-(4.96*10^5)log2/log.25=-2.48*10^5
..
half-life of material=2.48*10^5 yrs